OpenStudy (alyssacheeseburger):

I need help solving: 2sin(2x)-2sinx+2sqrt3cosx-sqrt3=0

OpenStudy (loser66):

sin(2x) = 2 sin x cos x, hence you have (4sin x cos x -2sin x)+ (2sqrt 3 cos x - sqrt 3)=0 2sin x (2 cos x -1) + sqrt 3 ( 2cos x -1) =0 (2cos x -1) (2 sin x + sqrt 3) =0 Now, easy, right?